How to integrate dv/DT=a-bv please tell fast my test is today

Grade:11

2 Answers

neptune1000
6 years ago
a= dv/dt,v=dr/dt.
hence, ∫dv=∫adt-b∫vdt=v-br+c=v
or, c=br. 
assuming of course a= acceleration,v=velocity,r= postion of particle
Arun
6 years ago
dv/dT = a-bv
dv/(a- bv) = dT
Now integrate it
- [ln(a- bv)]/b = T + constant
 
NOTE- you can also take constant in the form of ln
Hope it helps